BrushScriptX-Italic PostScript^{∗} Type-1 Fonts

Maurizio Loreti

Rev. 2.0 September 21, 2001

The files included in this package allow to use in your L^{A}TEX documents a PostScript Type-1 font named “BrushScriptX-Italic”; this font, similar to Adobe BrushScript, simulates hand-written characters. In detail, these files are:

1. AAA_readme.tex: this file.

2. README: a shorter ASCII version of this file.

3. Makefile: contains the commands needed, on my Unix system, to generate all the auxiliary files needed in order to use this fonts with L^{A}TEX and dvips. In particular, the included file generate.tex is used as input to the TEX package fontinst.

4. config.pbsi and pbsi.map: the auxiliary files for dvips. pbsi.map maps

BrushScriptX-Italic to the ASCII version of the font (the included file

BrushScriptX-Italic.pfa).

5. t1pbsi.fd: the TEX font definition file, for the T1 encoding.

6. pbsi8t.vf: the TEX virtual font file, for the T1 encoding.

7. pbsi8t.tfm, pbsi.tfm and pbsi8r.tfm: the TEX font metric files.
8. BrushScriptX-Italic.pfa and BrushScriptX-Italic.afm: the font, in

ASCII format, and the corresponding Adobe Font Metric file. The .afm file lacks the part about the kerning, that has been separately tuned by hand and is defined in another file, called kern.txt (merged with

BrushScriptX-Italic.afm at the generation time from Makefile).

9. sample.tex: a first example file (a real page printed out in the past for my students; in Italian, sorry).

10. pbsi.sty and example.tex: the style file defines a command \textbsi, that print its argument using the BrushScriptX-Italic fonts; the second file is another (dumb) example showing how to use that command.

∗PostScript is a registered trademark of Adobe Systems Inc.

For the installation, you should copy BrushScriptX-Italic.pfa, pbsi.sty, config.pbsi, pbsi.map, and all the the .fd, .tfm and .vf files, in appropriate directories of the TEX tree. Some installations (e.g. teTEX) require the update of the files database, that may be performed running (as root) the command texhash; check your documentation. Done that, use the fonts as shown in the sample.tex example file (for long texts), or (as outlined in example.tex) with \textbsi{...} (for short texts); and remember to run dvips with the command option -P pbsi. Some historical notes:

• I have found the original BrushScriptX-Italic font (a public domain NeXT screen font) on the web; and I have heavily modified the original files on my Amiga personal computer, with a commercial font drawing package called TypeSmith. This happened in the first years of 1990’s.

• I then submitted to CTAN the modified font; my version was also public domain, and still is: you are free to modify my files as you want, as long as you don’t make any profit from them.

• Later, when a new version of dvips appeared that made more tight controls on the handled Type-1 fonts, the BrushScriptX was refused by dvips. At the time I did not have any Amiga computer, nor any tool to use in order to edit and/or modify a Type-1 font; Louis Vosloo of Y&Y has pointed out to me that the original tool (TypeSmith) did not generate a correct font, and (with enormous courtesy) fixed the font himself. That modified font was resubmitted to CTAN.

• Still later, Mr. Rolf Niepraschk of Physikalisch-Technische Bundesanstalt Berlin (Germany) has both contributed a first version of the file pbsi.sty and, in 2001, found that the last dvips (5.86e) still refused to handle BrushScriptX. He has also contributed a fix for that, using the “type1fix” script part of the TEXtrace package by Péter Szabó (available under the

URL http://www.inf.bme.hu/~pts/textrace/).

• A kind reader of the comp.text.tex newsgroup pointed to me the existence

of a free font editor for Unix (http://pfaedit.sourceforge.net); so I

decided to clean up with it the font shapes, to specify better kernings and to build a brand new release (v2.0, released at the end of September 2001). Mr. Thierry Bouche gave invaluable help with fontinst, whose documentation was, err, quite obscure to me.

Be careful: do not use TEX commands that use math characters, or that change the typeface (like \bfseries or \sffamily); BrushScriptX-Italic has no defined variants!

Thanks to:

• Louis Vosloo <support@yandy.com>, who helped with the font.

• Rolf Niepraschk <niepraschk@ptb.de>, who helped with the font.

• Thierry Bouche <bouche@mozart.ujf-grenoble.Fr>, who helped with
fontinst.

• George Williams <gww@silcom.com>, the author of pfaedit. Happy TEXing! Maurizio Loreti

                                       EMail: loreti@pd.infn.it
                  WWW: http://www.pd.infn.it/~loreti/mlo.html